OpenStudy (anonymous):

help me with this question....... re- arrange the formulae in terms of x.... 2x-9=17

OpenStudy (anonymous):

2x - 9 = 17 2x = 17 + 9 2x = 26 x = 26/2 x = 13

OpenStudy (anonymous):

okay thanks........ here is a next one...... 3x-7=4

OpenStudy (anonymous):

Lol try and figure out the next one on your own, otherwise you won't learn

OpenStudy (anonymous):

gosh

OpenStudy (anonymous):

i dnt kno how to o this one thats why im asking you ........

OpenStudy (anonymous):

but never mind :(

OpenStudy (anonymous):

ill help you haha

OpenStudy (anonymous):

3x-7=4 3x = 4 +7 x = 11 x = 11/3

OpenStudy (anonymous):

so 11/3 is the answer

OpenStudy (anonymous):

yaaa

OpenStudy (anonymous):

okay thanks...... find the vallue of x.... 2x=7+4

OpenStudy (anonymous):

i dnt understand that

OpenStudy (anonymous):

i cant help anymore sorry lol, big test tomorrow $:

OpenStudy (anonymous):

gosh man

OpenStudy (anonymous):

dont stress, bump this, someone will help you

OpenStudy (anonymous):

okay thanks u anyways

OpenStudy (anonymous):

no probz!